OpenStudy (meowlover17):

EASY QUESTION JUST MAKING SURE QUESTION IS IN COMMENTS will give medal and fan :D

OpenStudy (meowlover17):

http://gyazo.com/3f99049c1489bb76a706c6f90fd00f44

OpenStudy (meowlover17):

I think the answer is A

imqwerty (imqwerty):

the answer is B (2^7)^(1/3) = 2^7/3 remember (a^b)^c = a^bc